ROTATIONAL SPECTRUM AND STRUCTURE OF Ar_{2}-H_{2}S COMPLEX: FT MICROWAVE SPECTRO-SCOPIC AND AB INITIO STUDIES

\underline{PANKAJ K. MANDAL} and E. ARUNAN, Department of Inorganic and Physical Chemistry, Indian Institute of Science, Bangalore- 560 012, INDIA.

Recently there has been a great deal of interest on the second row hydrides and their weakly bonded complexes and comparison to the first row hydrides. Complexes formed by H_{2}O and H_{2}S offer a wealth of information about intermolecular interaction involving first/second row hydrides. Most of the H_{2}O and H_{2}S complexes give at least two sets of rotational transitions corresponding to different internal rotor/tunneling states. In the case of Ar-H_{2}S complex, an anomalous substitution effect was observed.^{a} The rotational constant for Ar-D_{2}S was larger than that of Ar-H_{2}S, which is counter-intuitive. For Ar_{3}-H_{2}S, the isotope substitution effect was normal.^{b} What would be the situation for Ar_{2}-H_{2}S? Rotational transitions (asymmetric top) were predicted on the basis of an assumed geometry (C_{2}v). A series of a dipole transitions have been observed for Ar_{2}-H_{2}^{32}S. The absence of transitions between the states whose Kp and Ko are oo and oe indicates the presence of C_{2} axis interchanging two identical Ar nuclei (I=0). Twenty-one transitions have been fitted to three rotational constants and five centrifugal distortion constants and these are: A=1733.0982(11), B=1617.6570(5), C=830.2755(2) MHz; d1=9.270(5), d2=25.58(2), D_{J}=22.20(1), D_{JK}=3.34(4) and D_{K}=6.17(8)kHz. The rms deviation of the fit is 2.3 kHz, which is comparable to the experimental resolution. Rotational constant A is very close to that of the free Ar_{2} and this is the case in most of the Ar_{2}-HX complexes. Assuming H_{2}S to be spherical, the rotational constants are fitted to give a Ar_{2}(c.m.) to H_{2}S(c.m.) distance of 3.62 Å, and this leads to an Ar! -H_{2}S distance of 4.09 Å, which is comparable to the Ar-H_{2}S distances found in Ar_{m}-H_{2}S for m = 1 and 3. One more set of transitions has been observed and it is probably due to the ground (I = 0) internal rotor state. Further studies on this state and other isotopomers are in progress. Results will be presented in the symposium.

aH. S. Gutowsky, T. I. Emilsson and E. Arunan, J. Chem. Phys.\underline{106}, 5309 (1997).

bE. Arunan, T. I. Emilsson and H. S. Gutowsky and C. E. Dykstra, J. Chem. Phys.\underline{114}, 1242 (2001).